### v1

# Chemotaxis-induced phase separation

Abstract
Chemotaxis allows single cells to self-organize at the population level, as classically described by Keller–Segel models. We show that chemotactic aggregation can be understood using a generalized Maxwell construction based on the balance of density fluxes and reactive turnover. This formulation implies that aggregates generically undergo coarsening, which is interrupted and reversed by cell growth and death. Together, both stable and spatiotemporally dynamic aggregates emerge. Our theory mechanistically links chemotactic self-organization to phase separation and reaction–diffusion patterns.
Chemotaxis is a prevalent strategy in biology that enables both microbes and single (tissue) cells to move collectively and directionally in response to chemical gradients [1, 2, 3, 4, 5]. This can lead to self-organized aggregation when cells respond to gradients of self-secreted signaling molecules. The basic model describing chemotactic aggregation was introduced by Keller and Segel [6]. Diverse model variants account for different chemotactic responses, crowding effects, and cell proliferation [7]. Without cell proliferation and death the chemotactic aggregates are observed to undergo a coarsening process [7, 8, 9, 10]. In contrast, cell growth and death induce a fixed pattern wavelength leading to intricate patterning [11, 7], and even spatiotemporal chaos [12, 13]. However, a unifying understanding of chemotactic aggregates and their long-time dynamics in general Keller–Segel models is missing. It remains unknown whether a generic mechanism links coarsening, the selection of fixed intrinsic wavelengths, and the spatiotemporally dynamic patterns.
In this letter, we derive the generic behavior of chemotactic aggregates in Keller–Segel models by developing a combined description of the counteracting effects of random (diffusive) and chemotactic (directed) cell movement. Stationary aggregates are thereby constructed in phase space based on a reactive area balance representing a generalized Maxwell construction. We then argue that the resulting net cell movement drives the coarsening of collections of well-separated aggregates. (Slow) cell growth and death interrupt and reverse the coarsening process, resulting in both fully nonlinear steady-state patterns with fixed wavelengths (aggregate sizes) and spatiotemporally dynamic aggregates. Our theory is independent of the specific mathematical form of the models and relates the macroscopic dynamics of chemotactic aggregates to (active) phase separation [14, 15, 16] and (mass-conserving) reaction–diffusion systems [17, 18].
We consider a chemotactic population $\rho(\mathbf{x},t)$ subject to diffusion, with diffusion coefficient $D_{\rho}$, and chemotaxis, modeled as advection along the gradients in the chemoattractant field $c(\mathbf{x},t)$ [6]:
$\partial_{t}\rho(\bm{x},t)=D_{\rho}\nabla^{2}\rho-T\nabla\big{[}\chi_{\rho}(\rho)\chi_{c}(c)\rho\nabla c\big{]}+\varepsilon s(\rho).$ (1)
Here, ${T\chi_{\rho}(\rho)\chi_{c}(c)>0}$ represents the strength of chemotaxis with the positive sensitivity functions $\chi_{\rho}$ and $\chi_{c}$. For example, $\chi_{\rho}$ may describe crowding effects that suppress chemotaxis, while $\chi_{c}$ may capture nonlinear kinetics between the chemoattractant and its receptor in the cells [7]. Cell growth and death are described by $s(\rho)$ with a dimensionless strength $\varepsilon$. The dynamics of the chemoattractant, produced by the cells and subject to degradation, is given by ${\partial_{t}c(\bm{x},t)=D_{c}\nabla^{2}c+f(\rho,c)}$. To ensure a net chemoattractant production by the cells, we assume ${\partial_{\rho}f>0}$.
Random cell motion and chemotaxis can be captured in one quantity by rewriting Eq. (1) as
$\partial_{t}\rho(\bm{x},t)=T\nabla\big{[}\chi_{\rho}\rho\nabla\eta\big{]}+\varepsilon s(\rho)\,,$ (2)
in terms of the mass-redistribution potential (cf. Refs. [19, 20, 21])
$\eta(\bm{x},t)=\frac{D_{\rho}}{T}\int\mathrm{d}\rho\,\frac{1}{\chi_{\rho}\rho}-\int\mathrm{d}c\,\chi_{c}\,.$ (3)
Comparing the fluxes in Eq. (2) with the chemotaxis term in Eq. (1), it is evident that $\eta(\mathbf{x},t)$ acts as an effective chemorepellent. Random cell motion, appearing as a separate diffusion term in Eq. (1), is included in the population dynamics, Eq. (2), as one of two terms in the effective chemotaxis field $\eta$. Thereby, diffusion is treated as a modification to the chemotactic response of the cells, and the effective field $\eta$ integrates the counteracting effects of diffusion and chemoattraction.
Mathematically, the mass-redistribution potential $\eta$ resembles the chemical potential in binary phase separation with the mobility $\chi_{\rho}\rho$ [cf. Eq. (2) for ${\varepsilon=0}$ with Model B dynamics] [22, 23]. However, the mass-redistribution potential is in general not the derivative of a free energy functional but follows its own dynamic equation, as detailed in the Supplemental Material [24]. For particular model choices one can find a relation to a mathematical variational structure [25, 26]. For numerical analysis, we specifically consider the minimal Keller–Segel (mKS) model with ${\chi_{\rho}=\chi_{c}=1}$ and a linear chemoattractant production and decay term ${f=\rho-c}$ [27]. Then, one has ${\eta=\frac{D_{\rho}}{T}\log\rho-c}$. We start our analysis with the mass-conserving dynamics where ${\varepsilon=0}$.
Figure 1: (a,b) Illustration of the lateral instability. The mass-redistribution potential (orange) induces redistribution (orange arrows) of the cell density (blue) that counteracts (a) or amplifies (b) the initial perturbation (black arrows). (c) A $\mathsf{\Lambda}$-shaped nullcline leads to stationary peak patterns. The low-density intersection between the nullcline (black) and the flux-balance subspace ${\eta=\eta_{\mathrm{stat}}}$ (orange) determines the low-density plateau $\rho_{-}$ (pattern minimum ${\check{\rho}\gtrsim\rho_{-}}$). The value $\eta_{\mathrm{stat}}$ is determined by the reactive area balance (balance of the red-shaded areas). (d) A $N$-shaped nullcline gives a third high-density NC-FBS intersection determining a high-density plateau $\rho_{+}$ (pattern maximum ${\hat{\rho}\lesssim\rho_{+}}$). The resulting pattern is mesa-shaped.
Lateral instability. — A uniform cell density is unstable if chemotactic drift dominates over diffusive spreading. This requires the mass-redistribution potential $\eta$ to decrease where $\rho$ increases, such that gradients in $\eta$ guide motion towards increased cell densities [Eq. (2); cf. orange arrows in Fig. 1(a,b)]. For long-wavelength modulations, mass redistribution is slow compared to local reactions $f$ causing $\eta$ to quickly relax locally to its quasi-steady-state (QSS) $\eta^{*}(\rho)$, given by the nullcline (NC) ${f(\rho,c(\rho,\eta^{*}))=0}$ [insets in Fig. 1(a,b)]. Thus, chemotaxis-induced accumulation destabilizes uniform densities (lateral instability) if
$\partial_{\rho}\eta^{*}(\rho)<0\,.$ (4)
This instability is based on the self-amplifying redistribution of cell mass, and resemble the mass-redistribution instability found in mass-conserving reaction–diffusion systems as well as spinodal decomposition [20, 18]. Because this is a long-wavelength instability, the nullcline-slope criterion, Eq. (4), is exact beyond the QSS approximation; see Ref. [28].
Stationary aggregate profiles. — The stationary, one-dimensional pattern profile $[\tilde{\rho}(x),\tilde{c}(x)]$ can be determined in the $(\rho,\eta)$-phase plane by generalizing local equilibria theory [29, 20]. In stationary state, the diffusive and chemotactic fluxes must balance, i.e., the total flux ${\sim\partial_{x}\eta}$ [cf. Eq. (2)] must vanish everywhere on a domain with no-flux or periodic boundary conditions. Thus, stationary patterns are constrained to the flux-balance subspace (FBS) ${\tilde{\eta}(x)=\eta_{\mathrm{stat}}}$ [Fig. 1(c,d)]. Given that the densities ${\rho,c>0}$ are bounded from below, the (well-separated) patterns are either mesa-shaped and composed of low- and high-density plateaus connected by interfaces [Fig. 1(d)] or peak patterns. For peaks, the maximum density does not saturate in a plateau [Fig. 1(c)]. These patterns are constructed by noting that the profile curvature $\partial_{x}^{2}\tilde{c}$ vanishes (approximately) in plateau regions and at inflection points of the pattern. The stationary chemoattractant equation implies that those regions correspond to FBS-NC intersections. In particular, the low- and high-density FBS-NC intersections $\rho_{\pm}$ of a N-shaped nullcline are approached by the minimum $\check{\rho}$ and maximum $\hat{\rho}$ pattern densities exponentially in the plateau widths [sharp-interface approximation discussed in Supplemental Material [24]; Fig. 1(d)]. Peak patterns occur if the third FBS-NC intersection $\rho_{+}$ does not exist ($\mathsf{\Lambda}$-shaped nullcline) or lies at higher densities (asymmetric N-shaped nullcline) than the total domain mass $M$ [blue-shaded in Fig. 1(c,d)], fixed by the initial conditions, allows for. For example, the mKS model has a $\mathsf{\Lambda}$-shaped nullcline allowing unconstrained cell-density growth while receptor-binding kinetics [30] or volume-filling effects [31, 32] lead to N-shaped nullclines, directly explaining their density saturation. A precise, mathematical criterion for the transition from peak to mesa patterns is given in Ref. [33].
The FBS position $\eta_{\mathrm{stat}}$ was introduced previously as an integration constant [34]. Here, increasing $\eta$ at fixed density $\rho$ corresponds to decreasing the chemoattractant density $c$ [Eq. (3)]. Therefore, chemoattractant is produced (${f>0}$) above and degraded (${f<0}$) below the nullcline in the $(\rho,\eta)$-phase plane (multistability of the reaction kinetics is discussed in the Supplemental Material [24]). Consequently, the stationary value $\eta_{\mathrm{stat}}$ of a pattern adjusts such that the total production and degradation of the chemoattractant is balanced. This balance of the total turnover is represented (qualitatively) by a balance of areas enclosed between FBS and NC analogously to a Maxwell construction [Fig. 1(c,d); see Supplemental Material [24]]. For example, in the mKS model, the total turnover balance reads ${0=\int_{\check{\rho}}^{\hat{\rho}}\mathrm{d}\rho\,[\eta_{\mathrm{stat}}-\eta^{*}(\rho)]/\rho}$.
Figure 2: Illustration of the mass-competition instability for the competition of two peaks (a) and two plateaus (b; coalescence of the peak with its mirror image). In both cases (a) and (b), the changes in the stationary mass-redistribution potentials $\delta\eta$ of the two elementary patterns with the peak mass (blue-shaded areas) and plateau lengths, lead to self-amplifying mass transport (orange). (c) In a large system, these instabilities induce a coarsening process of the pattern as shown in two simulations with $T\bar{\rho}/D_{\rho}=1.5,3$. The faster mass-competition process [(inset in (d)] dominates the dynamics. (d) During the coarsening process, the average pattern wavelength $\langle\Lambda\rangle(t)$ continuously grows. Numerical simulations at $T\bar{\rho}/D_{\rho}=1.25,1.5,2,3$ (average over 3,1,8 and 2 replicates; light green to dark blue) are compared with the prediction ${\langle\Lambda\rangle(t)\sim\max(\gamma_{\pm})\log(t)}$ with ${\gamma\propto 5/8,3/4,1,1}$ (black lines; see Supplemental Material [24]). Apart from the average density $\bar{\rho}$ specified above, the simulation parameters are $T=5$, $D_{\rho}=0.1$, $D_{c}=1$, with system length $L=1000$.
Coarsening due to mass competition. — To understand the dynamics of patterns with several domains (peaks or mesas), we describe the interaction between two elementary patterns (two half-peak or two half-mesa patterns). As illustrated for peaks in the following, the pattern domains can interact either via competition for mass or coalescence with neighbors (mirror image) [Fig. 2(a,b); reflective boundaries]. Assuming a slow evolution of the pattern after each single elementary pattern relaxed to its (quasi-)stationary profile, the mass-redistribution potential at the single interfaces is given by its stationary value $\eta_{\mathrm{stat}}(M_{1,2})$ (QSS approximation). During competition [Fig. 2(a)], the dependence of $\eta_{\mathrm{stat}}(M_{1,2})$ on the peak masses ${M_{1,2}=\bar{M}\pm\delta M}$ induces a gradient $\eta(x)$ that redistributes mass between the peaks. The QSS value $\eta_{\mathrm{stat}}$ of the elementary pattern also depends on the plateau widths. Thus, during coalescence [Fig. 2(b)] the difference in the plateau widths induces a gradient $\eta(x)$ within the peak that induces peak motion.
The mass and plateau-length dependence of $\eta_{\mathrm{stat}}(M)$ arises because the reactive area balance depends on $\check{\rho},\hat{\rho}$, which vary with the peak mass and plateau lengths [Fig. 2(a,b)]. Associating changes in the low-density plateau lengths $\delta l$ with mass changes ${\delta M\approx\pm(\hat{\rho}-\check{\rho})\delta l}$, one describes both effects in terms of the mass dependence of $\eta_{\mathrm{stat}}$. To separate the effects of the peak height and the low-density plateau length, we denote the mass-induced change of the mass-redistribution potential only due to the change in $\hat{\rho}$ or $\check{\rho}$ by $\partial_{M}^{\pm}\eta_{\mathrm{stat}}$, respectively (see Ref. [28]). Importantly, because the reactive area balance does not depend on the specific mathematical model chosen, it generically implies ${\partial_{M}^{\pm}\eta_{\mathrm{stat}}<0}$ for both peak and mesa patterns; assuming for peak patterns that their height $\hat{\rho}$ grows with the mass $M$. Thus, mass is generically redistributed to the elementary pattern with initially higher mass, inducing a mass-competition instability. For peak competition [Fig. 2(a)], this leads to the collapse of the smaller and growth of the larger peak. For plateau competition [Fig. 2(b)], the shorter plateau vanishes due to the coalescence of the peak with its neighbor (mirror image). The growth rates of these two instabilities ${\sigma^{\pm}\sim-\partial_{M}^{\pm}\eta_{\mathrm{stat}}}$ depend on the strength of the $\eta$ gradient induced by the mass differences, i.e., the magnitude of $\partial_{M}^{\pm}\eta_{\mathrm{stat}}$ (see Supplemental Material [24]). Both competition processes drive the coarsening of patterns with several peaks or mesas until only one aggregate remains. Thus, uninterrupted coarsening is generic in mass-conserving Keller–Segel systems analogously to Ostwald ripening and coalescence in phase-separating systems and mass competition in reaction–diffusion systems [23, 17, 18].
In the mKS model, we find ${\sigma^{\pm}\sim\exp(-\Lambda/\gamma_{\pm})}$, $\sigma^{-}$ agreeing with the mathematical result [10]. Because one has ${\gamma_{-}\gtrless\gamma_{+}}$ for ${\bar{\rho}\gtrless 2D_{\rho}/T}$ [Fig. 2(d), inset], coalescence or competition is faster, respectively, and dominates the coarsening process at large wavelengths, which is verified numerically [Fig. 2(c)]. As ${[\sigma^{\pm}(\langle\Lambda\rangle)]^{-1}}$ is the typical timescale on which peaks vanish via collapse or coalescence at the mean pattern wavelength $\langle\Lambda\rangle$, the time during coarsening scales as ${t^{-1}\sim\sigma^{\pm}(\langle\Lambda\rangle)}$ [35, 36, 18]; see Supplemental Material [24]. The resulting coarsening law ${\langle\Lambda\rangle\sim\max(\gamma_{\pm})\log t}$ describes numerical simulations well, with deviations at the transition from competition- to coalescence-dominated coarsening [Fig. 2(d)].
Figure 3: (a) Illustration of interrupted coarsening due to production and degradation [$s(\rho)$; purple] counteracting the mass-competition instability (orange). (b) Plateau splitting occurs when the plateau density enters the laterally unstable region (green) due to production in the low-density plateau. In mesa patterns, high-density plateaus can split as well due to degradation when the plateau density decreases into the laterally unstable density regime. (c) The stability of patterns of the wavelength $\Lambda$ in the mKS model with a logistic source term ${s(\rho)=\rho(1-\rho)}$ of strength $\varepsilon$. Due to self-amplifying production in the low-density plateau ${s(\rho)\sim\rho}$, the threshold of plateau splitting (green) lies at low wavelengths $\Lambda$ and deforms the threshold of interrupted coarsening at large source strength $\varepsilon$ (solid compared to dashed blue line). The analytic thresholds (lines) well describe the numerical thresholds (circles and squares) obtained by numerical continuation and linear stability analysis (see Supplemental Material [24]). (d) The kymographs show $\log(1+\rho)$ (grayscale, 00 to $1$) for $\varepsilon=0.01,0.1$ simulated on a domain with reflective boundaries and length $L=149,74$ corresponding (approximately) to a multiple of the mean of the numerical thresholds $\Lambda_{\mathrm{stop,split}}$. In panels c and d, the same simulation parameters are used as in Fig. 2(d).
Source terms interrupt coarsening. — To account for cell growth and death, we now discuss the limit of weakly broken mass conservation (${\varepsilon\ll 1}$). Here, the average total density $\bar{\rho}_{\mathrm{stat}}$ of the stationary pattern is determined by the balance of overall production and degradation, ${\int_{\Omega}\mathrm{d}x\,s(\tilde{\rho})=0}$ (source balance), not by the initial condition. This condition requires ${0=s(\rho_{\mathrm{hss}})}$ at a density ${\check{\rho}<\rho_{\mathrm{hss}}<\hat{\rho}}$. Moreover, for stability of the elementary (half-mesa/-peak) pattern, the source term $s$ must drive any deviations $\bar{\rho}-\bar{\rho}_{\mathrm{stat}}$ from the average total density that balances production and degradation back to zero. We therefore assume net cell growth at low and net death at high densities as, for example, in logistic growth [cf. Fig. 3(a)]. As mass competition amplifies mass differences, and hence deviations in $\bar{\rho}$, a source term as specified above counteracts the mass-competition instability.
Consider the coalescence mode of a peak pattern [Figs. 2(b), 3(a)], the dominant coarsening pathway at large $\bar{\rho}_{\mathrm{stat}}$ in the mKS model [Fig. 2(d)]. If the peak shifts right by $\delta x$, the right plateau shortens, reducing the total production of cells by approximately ${-\varepsilon s(\rho_{-})\delta x}$ while production increases by the same amount to the left of the peak, as follows from integration of Eq. (2). Given that a peak shift $\delta x$ requires the addition/subtraction of mass $(\hat{\rho}-\rho_{-})\delta x$ to the right/left of the peak, the reduced production right and degradation to the left of the peak shifts the peak by ${(\hat{\rho}-\rho_{-})\partial_{t}^{\mathrm{source}}\delta x\approx-\varepsilon s(\rho_{-})\delta x}$. To this source-induced shift process ${\partial_{t}^{\mathrm{source}}\delta x}$ the competition process must be added. Together, one obtains the total mass-competition rate (for ${\varepsilon\ll 1}$; cf. Refs. [37, 17, 18] for the analogous effect in two-component reaction–diffusion systems)
$\sigma_{\varepsilon}^{-}(\Lambda)\approx\sigma_{\mathrm{D}}^{-}(\Lambda)-\varepsilon\frac{s(\rho_{-})}{\hat{\rho}(\Lambda)-\rho_{-}}.$ (5)
Here, $\sigma_{\mathrm{D}}^{-}(\Lambda)$ denotes the diffusion-limited part of the growth rate $\sigma^{-}$ that is dominant at large peak distances and masses (see Supplemental Material [24]). Thus, the increased source-induced mass influx to the peak from the left compared to the right counteracts the mass transport through the peak caused by the gradients in $\eta$ [Fig. 3(a)]. Mass competition is suppressed if ${\sigma_{\varepsilon}^{-}(\Lambda)\leq 0}$. Because the mass-competition rate ${\sigma_{\varepsilon}^{-}(\Lambda)}$ typically decreases strongly in $\Lambda$ (exponentially in the mKS model), the mass-competition instability is suppressed above a threshold wavelength $\Lambda_{\mathrm{stop}}$ fulfilling ${\sigma_{\varepsilon}^{-}(\Lambda_{\mathrm{stop}})=0}$ [dotted line in Fig. 3(c)]. Similar formulas hold for the competition mode and mesa patterns.
Plateau splitting at larger wavelengths. — As the source strength or the peak distance is increased further, the source term will increasingly deform the stationary profile compared to the pattern of the mass-conserving system because the production or degradation bends the plateaus inwards toward $\rho_{\mathrm{hss}}$ [Fig. 3(b)]. When the maximum of the low-density plateau (or the minimum of a high-density plateau) enters the region of negative nullcline slope ${\partial_{\rho}\eta^{*}<0}$ at wavelengths ${\Lambda\geq\Lambda_{\mathrm{split}}}$, the plateau loses stability, and a new peak or mesa is nucleated [Fig. 3(b)] [38, 39, 17]. Thus, splitting and interrupted coarsening discussed above steer the pattern toward a stationary, periodic configuration with a wavelength $\Lambda$ fulfilling ${\Lambda_{\mathrm{stop}}<\Lambda<\Lambda_{\mathrm{split}}}$ [Fig. 3(c)].
Successive coalescence and splitting. — The scale separation ${\Lambda_{\mathrm{stop}}\ll\Lambda_{\mathrm{split}}}$ can be lost if the source term (approximately) vanishes at a plateau density $\rho_{\pm}$, as in the mKS model with a logistic growth term ${s(\rho)=\rho(p-\rho)}$. Far below the splitting threshold $\Lambda_{\mathrm{split}}$, coalescence is counteracted by production ${\varepsilon s(\rho_{-})\approx\varepsilon\rho_{-}}$ in the low-density plateau ${\rho_{-}\approx 0}$ [Eq. (5); dotted line in Fig. 3(c)]. In contrast, near the splitting threshold, the plateau deformation inducing splitting is driven by a source term ${\sim\varepsilon\rho\gg\varepsilon\rho_{-}}$. Effectively, the source term appears weak for interrupted coarsening but strong for splitting, causing both thresholds to converge already at small source strength $\varepsilon$. Taking into account that the plateau density grows as the splitting threshold $\Lambda_{\mathrm{split}}$ is approached (see Supplemental Material [24]), the threshold of interrupted coarsening $\Lambda_{\mathrm{stop}}$ bends downward [solid blue line in Fig. 3(c)]. Nonetheless, the difference $\Lambda_{\mathrm{split}}-\Lambda_{\mathrm{stop}}$ decreases strongly with $\varepsilon$, well describing the numerical results in the mKS model [Fig. 3(c)].
As a result, at moderate source strength $\varepsilon$, several consecutive coalescence and splitting events occur before the stationary, periodic pattern develops [Fig. 3(d) top]. At larger source strength $\varepsilon$, the domain size would have to be fine-tuned for stationary periodic patterns to exist. Instead, the pattern continuously undergoes coalescence and splitting, reaching a dynamic steady state that has been argued to be chaotic [Fig. 3(d) bottom] [12].
Conclusions and outlook. —
Taken together, we provide a unified framework for understanding nonlinear patterns and their dynamics in chemotactic systems by introducing the mass-redistribution potential $\eta$. The gradients of this potential drive the migration of the chemotactic species integrating the counteracting effects of random and chemotactic cell movement. We have identified a geometric construction principle, akin to a Maxwell construction, that exploits flux-balance and reactive area balance. Thereby, our construction has a clear connection to the physics of the underlying chemotaxis system. As equilibrium-like pattern constructions were also obtained for different systems from the broad class of scalar active matter [40, 41, 42, 20, 43, 17, 44, 45, 46, 47], it is an exciting question how these constructions can be used to classify the coarse-grained dynamics of scalar active matter more generally.
Keller–Segel-type drift terms also arise if cells modulate motility based on quorum sensing [48, 49]. If instantaneous and (quasi) local this sensing leads to motility-induced phase separation. Also cross-diffusion gives rise to similar drift terms and phase-separation-like behavior [21]. Our results show more broadly that the coarse-grained dynamics of chemotactic systems resembles the phenomenology of phase separation.
Providing a unified understanding of generalized Keller–Segel models, we expect that our theory will provide valuable insights into multi-species chemotactic systems [26, 50, 51], the interplay with different aggregation effects like phase separation [52, 53, 54], and systems with more complex signaling [55]. In particular, it remains an intriguing question whether traveling fronts observed in systems with chemoattractant consumption, rather than production, or nutrient uptake can be analyzed using a similar framework [56, 57, 58, 59]. Moreover, understanding the effect of noise on wavelength selection and the spatiotemporally chaotic steady-state will be an exciting future challenge.
Acknowledgements.
We thank Fridtjof Brauns, Natan Dominko Kobilica, and Florian Raßhofer for inspiring discussions. We thank Alexander Ziepke for his critical reading of the manuscript. This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the Excellence Cluster ORIGINS under Germany’s Excellence Strategy – EXC-2094 – 390783311, the European Union (ERC, CellGeom, project number 101097810), and the Chan-Zuckerberg Initiative (CZI).